OpenStudy (callisto):

Problem of mathematical induction (MI) Prove by MI that (n+1)(n+2)(n+3)...(2n) = (2^n) . 1 .3 . 5 ... (2n-1)

OpenStudy (anonymous):

I would prefer a more readable version.

OpenStudy (callisto):

Let p(n) be the statement (n+1)(n+2)(n+3) ‧‧‧ (2n) = (2^n) ‧1 ‧ 3 ‧ 5 ‧‧‧ (2n-1)\] When n =1 LS = 1+1 =2 RS = (2^1)(2-1) = 2 So, p(1) is true. Assume p(k) is true When n =k+1 .... Then I can't continue :S

OpenStudy (callisto):

Or it should be I don't understand... Assume p(k) is true When n =k+1 LS = (k+2)(k+3) ‧‧‧ (2k)(2k+1)(2k+2) <- I don't understand the last 3 terms ...

OpenStudy (anonymous):

If p(k) is true, \[(k+1)\cdot(k+2)\cdots(2k) =2^k\cdot 1\cdots(2k-1)\] p(k+1) is,\[(k+2)\cdot(k+3)\cdots(2k+2)=2^{k+1}\cdot 1\cdots(2k)\]\[\implies (k+2)\cdots(2k)\cdot(2k+1)\cdot(2k+2)\cdot\frac{(k+1)}{(k+1)} = 2\cdot 2^k\cdot1\cdot2\cdots (2k-1)\cdot 2k\]

OpenStudy (callisto):

:| I think I got it now 2k = n-2 = 2(k+1) -2 = 2k 2k+1 = n-1 = 2(k+1) -1 = 2k+1 2k+2 = n = 2(k+1) = 2k+2 Right?!

OpenStudy (anonymous):

Typo, it's supposed to be\[\frac{(k+1)}{(k+1)}\cdot(k+2)\cdots(2k+2)=2\cdot2^{k}\cdot1\cdots(2k+1) \] Using p(k) on LHS. \[\text{LHS}=\frac{2(2k+1)(k+1)}{k+1}\cdot2^k\cdot1\cdots(2k-1)\]

OpenStudy (anonymous):

I might have missed something? Yes, I missed out the term 2k

OpenStudy (anonymous):

Rest is fine.

OpenStudy (anonymous):

Is it readable (my solution)? Do you understand it?

OpenStudy (callisto):

Your solution is readable. But you haven't answered my stupid question :|

OpenStudy (anonymous):

Yes, you did right.

OpenStudy (callisto):

Actually I did it in this way.. LS = (k+2)(k+3) ‧‧‧ (2k)(2k+1)(2k+2) = [2(k+1)(k+2)(k+3) ‧‧‧ 2k] (2k+1) = 2 ‧ (2^k) ‧ 1 ‧ 3 ‧ 5 ‧‧‧ (2k-1)(2k+1) = 2^(k+1) ‧ 1 ‧ 3 ‧ 5 ‧‧‧ (2k-1)([2(k+1)-1] = RS

OpenStudy (callisto):

Thanks!!!~ :)